# Calibrating doubly-robust estimators with unbalanced treatment assignment

### v1

@@ 1 Introduction
In this paper, we propose a simple extension of the DML estimator that addresses the issue of unbalanced treatment assignment. Inspired by the ML classification literature (Japkowicz and Stephen, 2002), the proposed approach undersamples the data used for fitting the ML model for the propensity score. Using the relation between the true and the undersampled propensity score, we calibrate the propensity scores to match the original data distribution. We show that the proposed estimator has the same asymptotic distribution as the DML estimator, attains the parametric rate of convergence $\sqrt{N}$ and its variance achieves the semi-parametric efficiency bound (Hahn, 1998). We illustrate the finite sample performance of the estimator in a simulation study.
@@ 2.4 Asymptotic results
The proof of Theorem 1 is relegated to the Appendix. Theorem 1 shows that the CU-DML estimator has the same asymptotic distribution as the DML estimator. In particular, the estimator attains the parametric rate of convergence $\sqrt{N}$ and its variance achieves the semi-parametric efficiency bound (Hahn, 1998).
@@ 3 Simulation study
In this section, we investigate the finite sample performance of the CU-DML estimator in a simulation study. We consider the following data generating process (DGP) used by Nie and Wager (2020):
@@
$Y_{i}$ $=b(X_{i})+(D_{i}-0.5)\left(X_{i,1}+X_{i,2}\right)+\sigma\epsilon_{i}.$
@@
The baseline main effect is the scaled Friedman (1991) function $b(X_{i})=\sin\left(\pi X_{i,1}X_{i,2}\right)+2\left(X_{i,3}-0.5\right)^{2}+X_{i,4}+0.5X_{i,5}$. For the propensity score we follow Künzel et al. (2019) and set $p(X_{i})=\alpha\left(1+\beta_{2,4}\left(\min(X_{i,1},X_{i,2})\right)\right)$ where $\beta_{2,4}(\cdot)$ is the beta cumulative distribution function with shape parameters 2 and 4. The share of treated is $\mathbb{E}[D_{i}]=(31/21)\alpha$. The innovation standard deviation is set to $\sigma\in\{1,5\}$. The following tuples of sample size and share of treated are considered: $\left(N,\mathbb{E}[D_{i}]\right)\in\{(2^{\prime}000,0.05),(2^{\prime}000,0.1),(4^{\prime}000,0.025),(4^{\prime}000,0.05),(4^{\prime}000,0.1),(8^{\prime}000,0.025),(8^{\prime}000,0.05),(8^{\prime}000,0.1)\}$. The case where $N=2^{\prime}000$ and $\mathbb{E}[D_{i}]=0.025$ is excluded since we would only have 50 treated observations. The simulation is repeated 1’000 times for each tuple and each $\sigma$ value. [^4] The code for the simulation study is availabile at https://github.com/dballinari/Simulation-CU-DML.
@@
We compare the CU-DML estimator with the DML estimator (Chernozhukov et al., 2018), and three popular adjustments of the DML estimator that are designed to improve its finite sample performance when the treatment assignment is unbalanced. [^5] Over the past two decades, these adjustments have been primarily employed to improve the finite sample performance of the inverse-probability-weighted estimator when the treatment assignment is unbalanced; see, among others, the extensive simulation study of Huber et al. (2013). More recently, these approaches have also been employed to improve the performance of the DML estimator. The W-DML estimator uses winsorized propensity scores at 0.01 and 0.99 (Imbens, 2004). In the N-DML estimator we normalize the weights $w_{i}^{\mathcal{I}_{-k}}:=D_{i}/\widehat{p}^{\mathcal{I}_{-k}}(X_{i})$ to sum to unity (e.g. Słoczyński and Wooldridge, 2018). Following Huber et al. (2013), the T-DML estimator normalizes the weights $w_{i}^{\mathcal{I}_{-k}}$ and then truncates them to not exceed 0.04, before normalizing them again. Finally, we also consider the U-DML estimator, where the entire sample is undersampled and the DML estimator is applied to the undersampled data. All estimators use 5-fold cross-fitting. The nuisance functions are estimated using random forests (Breiman, 2001) with 500 trees. The maximal depth of the trees and the minimal number of observations in the leaf nodes are determined by 5-fold cross-validation over 20 simulation replications and set to the most frequently selected values. [^6] Properely tuning the hyperparameters of the random forests is crucial, especially for DML when the treatment assignment is unbalanced. The popular Python implementation of random forest, scikit-learn (Pedregosa et al., 2011), sets the default minimal number of observations in the leaf nodes to 1. This may lead to highly unstable propensity scores when the treatment assignment is unbalanced. See Bach et al. (2024) for a recent simulation study on the importance of hyperparameter-tuning for the DML estimator.
Table 1: Statistics of the simulations
@@
  • Note: The table reports the root mean squared error (RMSE), the absolute value of the average bias, the standard deviation, and the coverage of the 95% confidence interval for the ATE estimators across 1’000 simulations. The best performance in terms of RMSE is highlighted in bold. The table reports the results for 16 different simulation settings, where the sample size $N$, the share of treated $\mathbb{E}[D_{i}]$, and the innovation standard deviation $\sigma$ are varied.
@@
The results are presented in Table 1. The table reports the root mean squared error (RMSE), the absolute value of the average bias, the standard deviation, and the coverage of the 95% confidence interval for each ATE estimator. The best performance in terms of RMSE in each simulation is highlighted in bold. The lowest RMSE is achieved by the CU-DML estimator in 10 out of 16 cases. Only when at least 400 observations are treated, the CU-DML estimator is outperformed in terms of RMSE by the normalized and trimmed DML estimators. However, the RMSE differences between N-DML, T-DML, and CU-DML estimators are considerably small in these cases. This pattern is in line with the findings from the machine learning classification literature, showing that class imbalancedness is a relative problem, related to both the degree of imbalancedness and the sample size (Japkowicz and Stephen, 2002).
@@
The bias and standard deviation of the estimators unveil that, while DML has the lowest bias among all estimators, CU-DML reduces the variance in the ATE estimator and has the lowest standard deviation across all estimators and settings. The simulation study also confirms the theoretical result presented in the previous section. First, the coverage is in all settings close to 95%. Second, the RMSE decreases at rate $\sqrt{N}$.

### v2

@@ 1 Introduction
In this paper, we propose a simple extension of the DML estimator that addresses the issue of unbalanced treatment assignment. Inspired by the ML classification literature (Japkowicz and Stephen, 2002), the proposed approach undersamples the data used for fitting the ML model for the propensity score. Using the relation between the true and the undersampled propensity score, we calibrate the propensity scores to match the original data distribution. We show that the proposed estimator has the same asymptotic distribution as the DML estimator, attains the parametric rate of convergence $\sqrt{N}$ and its variance achieves the semi-parametric efficiency bound (Hahn, 1998). We illustrate the finite sample performance of the estimator in a simulation study. While we present the results in the context of the DML estimator, the proposed approach applies to any ATE estimator that relies on the efficient score function.
@@ 2.4 Asymptotic results
The proof of Theorem 1 is relegated to Appendix A. Theorem 1 shows that the CU-DML estimator has the same asymptotic distribution as the DML estimator. In particular, the estimator attains the parametric rate of convergence $\sqrt{N}$ and its variance achieves the semi-parametric efficiency bound (Hahn, 1998).
@@ 3 Simulation study
In this section, we investigate the finite sample performance of the CU-DML estimator in a simulation study. We consider two simulation strategies. In the first strategy, we generate data from a synthetic data generating process (DGP). In more detail, we generate data from the DGP used by Nie and Wager (2020):
@@
$Y_{i}$ $=b(X_{i})+(D_{i}-0.5)\left(X_{i,1}+X_{i,2}\right)+\sigma\epsilon_{i}.$ (8)
@@
The baseline main effect is the scaled Friedman (1991) function $b(X_{i})=\sin\left(\pi X_{i,1}X_{i,2}\right)+2\left(X_{i,3}-0.5\right)^{2}+X_{i,4}+0.5X_{i,5}$. For the propensity score we follow Künzel et al. (2019) and set $p(X_{i})=\alpha\left(1+\beta_{2,4}\left(\min(X_{i,1},X_{i,2})\right)\right)$ where $\beta_{2,4}(\cdot)$ is the beta cumulative distribution function with shape parameters 2 and 4. The share of treated is $\mathbb{E}[D_{i}]=(31/21)\alpha$ and the true ATE is 1. The innovation standard deviation is set to $\sigma=1$, and results for $\sigma=5$ are relegated to Appendix B.
In the second strategy, we analyze the CU-DML estimator’s performance using an Empirical Monte Carlo Study (EMCS) approach (Huber et al., 2013; Lechner and Wunsch, 2013). Our EMCS, following Knaus et al. (2022), uses a dataset from Lechner et al. (2020) of Swiss unemployed individuals in 2003. The dataset includes data on the effect of a job search program (treatment) on the number of months employed in the first six months after the start of the program (outcome), and 49 covariates providing information on individual characteristics of the unemployed, the regional employment agency, and regional labour market characteristics. For more data details, see Knaus et al. (2022). The EMCS proceeds as follows. We estimate the propensity score using a logistic regression on the entire sample of 91’339 unemployed individuals. Then, we define a new sample consisting of only the non-treated individuals whose fitted propensity score $\hat{p}_{i}$ lies between $0.05$ and $0.95$. In each simulation, we draw a random sample with replacement of size $N$ from this new sample and randomly assign treatments as $D_{i}=\mathds{1}{\{V_{i}<\hat{p}_{i}/\lambda\}}$, where $\hat{p}_{i}$ is the fitted propensity score from the logistic regression, $V_{i}\sim\text{Unif}[0,1]$ and $\lambda$ controls the share of treated $\mathbb{E}[D_{i}]$. The true ATE is 0 by construction.
@@
We compare the CU-DML estimator with the DML estimator (Chernozhukov et al., 2018), and three popular adjustments of the DML estimator that are designed to improve its finite sample performance when the treatment assignment is unbalanced. [^4] Over the past two decades, these adjustments have been primarily employed to improve the finite sample performance of the inverse-probability-weighted estimator when the treatment assignment is unbalanced; see, among others, the extensive simulation study of Huber et al. (2013). More recently, these approaches have also been employed to improve the performance of the DML estimator. The W-DML estimator uses winsorized propensity scores at 0.01 and 0.99 (Imbens, 2004). In the N-DML estimator we normalize the weights $w_{i}^{\mathcal{I}_{-k}}:=D_{i}/\widehat{p}^{\mathcal{I}_{-k}}(X_{i})$ to sum to unity (e.g. Słoczyński and Wooldridge, 2018). Following Huber et al. (2013), the T-DML estimator normalizes the weights $w_{i}^{\mathcal{I}_{-k}}$ and then truncates them to not exceed 0.04, before normalizing them again. Finally, we also consider the U-DML estimator, where the entire sample is undersampled and the DML estimator is applied to the undersampled data. All estimators use 5-fold cross-fitting. The nuisance functions are estimated using random forests (Breiman, 2001) with 500 trees. The maximal depth of the trees and the minimal number of observations in the leaf nodes are determined by 5-fold cross-validation over 20 simulation replications and set to the most frequently selected values. [^5] Properely tuning the hyperparameters of the random forests is crucial, especially for DML when the treatment assignment is unbalanced. The popular Python implementation of random forest, scikit-learn (Pedregosa et al., 2011), sets the default minimal number of observations in the leaf nodes to 1. This may lead to highly unstable propensity scores when the treatment assignment is unbalanced. See Bach et al. (2024) for a recent simulation study on the importance of hyperparameter-tuning for the DML estimator. Details on the software used for the simulations are provided in Appendix B.1.
Table 1: Statistics of the two simulation strategies
Panel A: Synthetic DGP
@@
Panel B: Empirical Monte Carlo Study
$N=2000$, $\mathbb{E}[D_{i}]=2.5\%$ $N=4000$, $\mathbb{E}[D_{i}]=2.5\%$ $N=8000$, $\mathbb{E}[D_{i}]=2.5\%$
@@
DML - - - - 0.314 0.051 0.309 0.984 0.194 0.041 0.189 0.975
U-DML - - - - 0.261 0.024 0.260 0.950 0.183 0.013 0.183 0.962
CU-DML - - - - 0.208 0.016 0.208 0.946 0.148 0.023 0.146 0.944
W-DML - - - - 0.233 0.044 0.228 0.979 0.162 0.051 0.154 0.961
N-DML - - - - 0.253 0.057 0.246 0.948 0.174 0.050 0.166 0.940
T-DML - - - - 0.237 0.051 0.231 0.937 0.167 0.052 0.159 0.938
$N=2000$, $\mathbb{E}[D_{i}]=5.0\%$ $N=4000$, $\mathbb{E}[D_{i}]=5.0\%$ $N=8000$, $\mathbb{E}[D_{i}]=5.0\%$
DML 0.256 0.049 0.252 0.977 0.172 0.043 0.167 0.962 0.111 0.032 0.107 0.959
U-DML 0.262 0.012 0.261 0.964 0.180 0.012 0.179 0.953 0.129 0.022 0.127 0.948
CU-DML 0.213 0.015 0.213 0.958 0.143 0.023 0.141 0.943 0.101 0.023 0.099 0.941
W-DML 0.244 0.051 0.239 0.976 0.164 0.044 0.158 0.962 0.108 0.033 0.103 0.957
N-DML 0.233 0.056 0.226 0.950 0.163 0.048 0.156 0.941 0.110 0.036 0.104 0.941
T-DML 0.226 0.056 0.219 0.946 0.161 0.048 0.154 0.941 0.110 0.036 0.104 0.941
$N=2000$, $\mathbb{E}[D_{i}]=10.0\%$ $N=4000$, $\mathbb{E}[D_{i}]=10.0\%$ $N=8000$, $\mathbb{E}[D_{i}]=10.0\%$
DML 0.156 0.033 0.153 0.962 0.107 0.024 0.105 0.961 0.072 0.014 0.071 0.960
U-DML 0.184 0.016 0.183 0.962 0.128 0.017 0.127 0.955 0.088 0.014 0.087 0.952
CU-DML 0.146 0.020 0.144 0.957 0.099 0.016 0.098 0.950 0.069 0.012 0.068 0.949
W-DML 0.156 0.033 0.152 0.962 0.106 0.023 0.104 0.961 0.072 0.014 0.071 0.960
N-DML 0.154 0.037 0.149 0.946 0.107 0.026 0.103 0.952 0.072 0.016 0.070 0.951
T-DML 0.153 0.038 0.149 0.945 0.106 0.025 0.103 0.952 0.072 0.016 0.070 0.951
@@
  • Note: The table reports the root mean squared error (RMSE), the absolute value of the average bias, the standard deviation, and the coverage of the 95% confidence interval for the ATE estimators across 1’000 simulations. Panel A reports the results for the synthetic DGP where data is generated according to Equation (8), while Panel B reports the results for the Empirical Monte Carlo Study. For each of the two simulation strategies, the table reports the results for 8 different simulation settings, where the sample size $N$ and the share of treated $\mathbb{E}[D_{i}]$ are varied. For the synthetic DGP, the innovation standard deviation $\sigma$ is set to 1, which corresponds to a signal-to-noise ratio of $\mathrm{Var}[y]/\sigma^{2}=1.3$. The best performance in terms of RMSE is highlighted in bold.
@@
The results are presented in Table 1. The table reports the root mean squared error (RMSE), the absolute value of the average bias, the standard deviation, and the coverage of the 95% confidence interval for each ATE estimator. The simulation is repeated 1’000 times for each of the following tuples of sample size and share of treated: $\left(N,\mathbb{E}[D_{i}]\right)\allowbreak\in\allowbreak\{(2^{\prime}000,0.05),(2^{\prime}000,0.1),(4^{\prime}000,0.025),(4^{\prime}000,0.05),(4^{\prime}000,0.1),(8^{\prime}000,0.025),(8^{\prime}000,0.05),(8^{\prime}000,0.1)\}$. The case where $N=2^{\prime}000$ and $\mathbb{E}[D_{i}]=0.025$ is excluded since we would only have 50 treated observations. The best performance in terms of RMSE in each simulation is highlighted in bold. For the synthetic DGP (Panel A), the lowest RMSE is achieved by the CU-DML estimator in 5 out of 8 settings. Only when at least 400 observations are treated, the CU-DML estimator is outperformed in terms of RMSE by the normalized and trimmed DML estimators. However, the RMSE differences between N-DML, T-DML, and CU-DML estimators are considerably small in these cases. These results are unaffected by a higher innovation standard deviation $\sigma=5$, see Table B1. For the EMCS (Panel B), CU-DML achieves the smallest RMSE for all combinations of $\left(N,\mathbb{E}[D_{i}]\right)$. Also for the EMCS, we observe that the outperformance of the CU-DML estimator decreases for larger samples and/or less imbalanced samples. These patterns are in line with the findings from the machine learning classification literature, showing that class imbalancedness is a relative problem, related to both the degree of imbalancedness and the sample size (Japkowicz and Stephen, 2002).
@@
The bias and standard deviation of the estimators unveil that, while the analyzed estimators do not differ substantially in terms of their bias, the CU-DML estimator has the lowest standard deviation across all simulation strategies and settings. The simulation study also confirms the theoretical result presented in the previous section. First, the coverage of the proposed estimator is in all strategies and settings close to 95%. Second, its RMSE decreases at rate $\sqrt{N}$.
@@ Appendix A Appendix: Proofs
Appendix B Appendix: Details on the simulation study
B.1 Implementation details
The simulation study is implemented in Python 3.12. The data generating process is implemented using the numpy 1.26.4 package. The machine learning models are estimated using the scikit-learn 1.4.0 package, using the acceleration extension scikit-learn-intelex 2024.4.0. The empirical data is processed using pandas 2.2.2. The code for the simulation study is availabile at https://github.com/dballinari/Calibrating-doubly-robust-estimators-with-unbalanced-treatment-assignment.
B.2 Additional simulation results
Table B1: Statistics of the synthetic DGP with $\sigma=5$
DML - - - - 0.885 0.025 0.884 0.983 0.448 -0.001 0.448 0.980
U-DML - - - - 0.761 0.097 0.755 0.948 0.523 0.065 0.519 0.954
CU-DML - - - - 0.588 0.056 0.585 0.944 0.402 0.019 0.401 0.963
W-DML - - - - 0.675 0.053 0.673 0.974 0.429 0.005 0.428 0.979
N-DML - - - - 0.698 0.046 0.696 0.932 0.420 0.008 0.420 0.948
T-DML - - - - 0.671 0.055 0.669 0.934 0.420 0.008 0.420 0.948
  • Note: The table reports the root mean squared error (RMSE), the absolute value of the average bias, the standard deviation, and the coverage of the 95% confidence interval for the ATE estimators across 1’000 simulations. The data is generated from the synthetic DGP according to Equation (8), with innovation standard deviation $\sigma=5$, which corresponds to a signal-to-noise ratio of $\mathrm{Var}[y]/\sigma^{2}=1.0$. The table reports the results for 8 different simulation settings, where the sample size $N$ and the share of treated $\mathbb{E}[D_{i}]$ are varied. The best performance in terms of RMSE is highlighted in bold.